A rubber ball was launched at a velocity of +5 m/s at an angle where its range is at a maximum. A target is located exactly 2.0 meters from where the ball was launched.

a) What is the velocity of the ball as soon as it hits the ground?

b) Will the rubber ball hit the target? If no, at what angle should the ball be launched for it to hit the target?
1
Expert's answer
2020-05-03T17:17:34-0400

(a) The velocity of the ball as soon as it hits the ground

"v=5\:\rm m/s"

(b) The range

"R_{\rm max}=\frac{v_0^2}{g}=\frac{5^2}{9.8}=2.5\:\rm m\neq2.0\:\rm m"

Hence, the rubber ball doesn't hit the target.


"R=\frac{v_0^2\sin2\theta}{g}=\frac{5^2\sin2\theta}{9.8}=2.0"

"\theta=26^{\circ}"